# S3E: A Mulit-Robot Multimodal Dataset for Collaborative SLAM

### v6

@@ Abstract
Note to Practitioners
The S3E dataset is a valuable asset for professionals in robotics and autonomous systems, offering a rich, multimodal set of data to refine Collaborative Simultaneous Localization and Mapping (C-SLAM) technologies. With synchronized streams from LiDAR, cameras, IMUs, and UWB across indoor and outdoor environments, it presents a diverse range of real-world challenges for testing and developing C-SLAM algorithms. This dataset is particularly beneficial for evaluating how robots perform in collaborative tasks, with its unique trajectory paradigms simulating various operational scenarios. Practitioners can use S3E to enhance their systems’ ability to handle dynamic environments, spatial constraints, and limited sensor overlap. However, working with S3E requires significant computational resources and expertise in C-SLAM. The complexity of the data may be daunting for those new to the field. Despite this, the S3E dataset is an excellent platform for pushing the boundaries of collaborative robotics, providing a rigorous benchmark for current technologies and a foundation for future developments.
Index Terms:
Collaborative Simultaneous Localization and Mapping, Multi-Agent Systems, Dataset, Trajectory Paradigms
@@ I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is pivotal for enabling multi-agent cooperation, thereby enhancing the robustness and efficiency of concurrent localization and mapping endeavors within a shared operational space [1, 2, 3, 4, 5, 6, 7]. Despite considerable progress in C-SLAM research during the past decade, issues of repeatability and benchmarking remain significant hurdles within the domain. The field is primarily constrained by two limitations. First, the complexity of C-SLAM systems, which integrate intricate software architectures and a variety of sensors, complicates the process of achieving perfect replication, thereby emphasizing the need for publicly available datasets to facilitate methodological evaluation. Second, a common method for evaluating C-SLAM solutions involves partitioning a single-agent SLAM dataset into several segments, each assigned to a distinct virtual agent [8]. However, this methodology is impractical for real-world multi-agent scenarios, given the uniform perspectives and lighting conditions that all virtual agents would encounter at intersection points.
@@
Furthermore, the trajectory design for multi-robot scenarios must adhere to two fundamental principles. The first principle is temporal and spatial diversity. The trajectories should be designed to accommodate the varying intra- and inter-loop closures at different times and locations, thus offering complementary vantage points for the participating robots. The second principle pertains to the communication constraints. In practical applications, collaborative robots are limited to sharing information within a certain proximity. Hence, maintaining a reasonable interaction distance between robots within these regions is essential. In this paper, we introduce four trajectory prototypes that conform to diverse intra- and inter-robot closure strategies, aiming to evaluate the generalizability of C-SLAM methodologies.
Within the SLAM research community, publicly available datasets are indispensable for accelerating the development of SLAM algorithms suitable for real-world deployment. This significance stems from two primary factors. First, the execution of specialized SLAM experiments requires substantial investment in sophisticated hardware and complex software, including calibration routines and ground truth generation systems; hence, the availability of datasets substantially aids in the algorithmic development process. Second, these datasets provide a fair and impartial benchmarking framework for comparative analysis and performance assessment. Over the past decade, an array of SLAM datasets has emerged, each contributing uniquely to the diversity of sensor modalities and the breadth of scenarios and motion types they encompass [9, 10]. Nonetheless, the preponderance of these datasets is skewed towards single-agent systems, equipped with an assortment of sensors such as cameras, inertial measurement units (IMUs), and LiDAR, which may limit their applicability to multi-agent contexts.
The limited availability of datasets specifically tailored for C-SLAM is largely due to the escalating challenges in data acquisition as the number of collaborating agents increases. Existing C-SLAM datasets, such as those from [11, 12, 13, 14, 15, 16], predominantly feature a single cooperative paradigm with substantial spatial overlap among the robots, often overlooking cooperative strategies in areas with minimal overlap. To address this gap and bolster ongoing research in C-SLAM, we introduce S3E, a comprehensive multimodal dataset. S3E captures a variety of cooperative trajectory patterns across both outdoor and indoor settings.
Utilizing our S3E dataset as a foundation, we have conducted an extensive series of experiments employing state-of-the-art SLAM methodologies. These experiments spanned the spectrum of SLAM applications, from single-robot to collaborative scenarios, incorporating a rich array of sensor data. Our results indicate that current collaborative SLAM techniques often encounter suboptimal performance, particularly in handling inter-loop closures across certain sequences. This finding highlights the critical need for ongoing research aimed at improving the robustness and reliability of C-SLAM systems.
@@
  • We have assembled a pioneering C-SLAM dataset through the deployment of three state-of-the-art ground robots. Each robot is equipped with an advanced 16-beam 3D laser scanner, two high-resolution color cameras, a 9-axis IMU, a UWB receiver, and a dual-antenna RTK receiver. Notably, this is the first C-SLAM dataset to include UWB relative distance measurements, offering a novel dimension for research.
  • To evaluate the adaptability of C-SLAM in environments with sparse overlap, we have captured extensive long-term sequences. These were obtained using four carefully crafted trajectory paradigms that account for a variety of intra- and inter-robot loop closure scenarios.
  • We have undertaken an in-depth evaluation of contemporary C-SLAM techniques and their single-agent counterparts. This assessment leverages a wide array of sensor data, with a meticulous focus on the unique characteristics and performance attributes of each method.
[…]
[autoplay,loop,width=]5videos/360/dormitory_1/000119
Figure 1: Animations of Dormitory_1. Data is captured simultaneously by a trio of Unmanned Ground Vehicles (UGVs). Each UGV initiates its journey from distinct starting points and independently explores separate areas. Upon successful detection of inter-loop closures, they exchange valuable information and collaboratively construct a globally consistent map. For optimal viewing of the animations, it is recommended to utilize Adobe Reader or KDE Okular.
@@ II Related Work
Collaborative SLAM. The integration of collaborative agents within the framework of SLAM presents a significant advancement, offering the potential for increased operational velocity and efficiency. This collaborative approach has the capacity to outstrip the performance of single-agent systems [17, 18]. LABEL:Fig:ani_dorm_1 demonstrates the operational concept of C-SLAM, where a collective of robots is enabled to explore various sectors of an environment in parallel, thereby enhancing the overall efficiency of the exploration process compared to traditional single-agent methodologies.
C-SLAM systems can be differentiated based on their underlying network topologies, which include centralized, distributed, and adaptive models. In a centralized topology, a central server is tasked with the aggregation of all incoming sensory data, ensuring a unified estimation of multiple trajectories, and subsequently distributing these findings to the constituent agents [19, 20, 2]. Conversely, in distributed topologies, computational responsibilities are decentralized among the team of agents, with some systems utilizing a message broadcasting protocol for the propagation of information [21]. However, scalability becomes a formidable challenge in such distributed frameworks, particularly in scenarios involving a large number of robots, due to limitations in bandwidth and communication range.
To mitigate these challenges, adaptive strategies have been proposed, advocating for the transmission of information within the immediate communication sphere of the agents [22, 23, 3]. Swarm-SLAM [24] represents an innovative paradigm within this domain, adept at minimizing the communication load between robots while simultaneously accelerating the convergence of the SLAM process through the implementation of sophisticated prioritization algorithms for inter-robot loop closures. Moreover, hybrid approaches present a viable alternative, wherein a subset of robots engages in data processing, while others are dedicated to the application of the outcomes derived from this processing [25].
@@
SLAM Datasets. High-quality and extensive datasets are indispensable for propelling advancements in SLAM research. Over recent years, there has been a commendable effort to release datasets to the research community. The KITTI dataset [9] is particularly noteworthy for its extensive and high-quality urban data, serving as a benchmark for both vision-based and LiDAR-related SLAM techniques. However, KITTI is not without limitations; it lacks perfectly synchronized data, offering only estimated time delays between different sensors. To surmount this limitation, the EuRoC dataset [10] was collected, providing synchronized visual and inertial data captured by a micro aerial vehicle in controlled environments, synchronized through the use of hardware clocks.
Despite these commendable efforts, there is a recognized scarcity of datasets tailored for multi-agent scenarios. The UTIAS dataset [11] stands as a pioneering work in multi-agent SLAM, featuring a collaborative system of five robots exploring a small area with identifiable landmarks. The AirMuseum dataset [13] introduces a heterogeneous multi-robot SLAM dataset within a larger warehouse environment, offering ground-truth trajectories computed using Structure-from-Motion (SfM) [26] and constrained with fixed beacons. The FordAV dataset [12] presents a dataset of multiple autonomous vehicles with a 3D map, capturing seasonal variations in dynamic urban environments. However, these datasets tend to have substantial overlap in the co-view region, which may limit the comprehensive evaluation of C-SLAM robustness.
In an effort to address the limitations of existing datasets, the GRACO dataset [15] presents a novel approach by introducing a ground-aerial dataset. This dataset leverages the drone’s altitude to dynamically adjust the extent of view overlap, offering a unique perspective on multi-agent SLAM scenarios. However, GRACO exhibits some inconsistencies in the synchronization of data acquisition between aerial and ground perspectives, and it does not fully accommodate the complexities of dynamic environments.
The SubT-MRS dataset [16] contributes to the field by providing a diverse array of multi-robot SLAM data captured across a variety of environments. This dataset enhances the representation of real-world applications and the challenges posed by different settings.
Kimera-Multi [14] further advances the state of C-SLAM datasets by offering a comprehensive collection characterized by its diverse trajectory designs and the involvement of multiple robots. This dataset is instrumental in studying the collaborative aspects of multi-agent systems.
@@
It is crucial to acknowledge, however, that existing datasets, including GRACO, SubT-MRS, and Kimera-Multi, often do not fully consider the intricacies of collaborative trajectory patterns. The robustness of C-SLAM systems is influenced by the ability to handle intra- and inter-loop closures, a factor that is not consistently addressed in current datasets. The consideration of these collaborative patterns is essential for developing C-SLAM algorithms that can effectively operate in environments with varying levels of agent interaction and dynamic scene elements.
@@
In this paper, we introduce a large-scale C-SLAM dataset that includes visual ambiguities and dynamic objects, captured under four distinct trajectory paradigms to assess C-SLAM generalizability under restricted overlap conditions. Our dataset also incorporates UWB relative measurement positioning, adding a significant dimension for C-SLAM research.
For a comprehensive comparison of different datasets, including our contribution, refer to LABEL:Tab:dataset_comparison.
Table II: Payload Sensor and Ground Truth Device Specifications. It includes information about the types of sensors, their respective capabilities, and the ground truth devices used for validation and accuracy assessment.
Device Type Unit
LiDAR Velodyne
   Model VLP-16 Puck
   Channels $16$
   Range $100$ $m$
   Accuracy ±$\pm 3$ $cm$
   Vertical FOV ±$\pm 15$
   Vertical Resolution $2$
   Horizontal FOV $360$
   Horizontal Resolution $0.1\sim 0.4$
   Max. Frequency $10$ $Hz$
Camera - Stereo HikRobot
   Model MV-CS050-10GC
   Data Interface GigE
   Sensor Model Sony IMX264
   Resolution $1224\times 1024$ $pixel$
   Dynamic Range $72$ $dB$
   Baseline $360$ $mm$
   Max. Frequency $24.2$ $Hz$
   Lens Model MVL-MF0828M-8MP
   Focal Length 8 $mm$
   Vertical FOV 54.97
   Horizontal FOV 68.46
IMU Xsens
   Model MTi-30-2A8G4
   Gyro In-Run Bias Stability $18$ ${}^{\circ}/h$
   Accel In-Run Bias Stability $15$ $\mu g$
   Mag RMS Noise 0.5 $mGauss$
   Max. Frequency 400 $Hz$
UWB NoopLoop
   Model LinkTrack P-B
   Max. Distance $500$ $m$
   Accuracy ±$\pm 10$ $cm$
   Max. Frequency $200$ $Hz$
RTK Femtomes
   Model Nano-D
   RTK Accuracy ±$\pm 1$ $cm$
   Max. Frequency 20 $Hz$
GNSS/INS CHCNAV
   Model CGI-610
   RTK Accuracy ±$\pm 1$ $cm$
   Gyro In-Run Bias Stability 2.5 ${}^{\circ}/h$
   Accel In-Run Bias Stability 15 $\mu g$
   Max. RTK Frequency 20 $Hz$
   Max. INS Frequency 100 $Hz$
Motion Capture Vicon
   Model Tracker 3.8
   Camera Model Vero 2.2
   Num. of Cameras 17
   Position Accuracy ±$\pm 1$ $mm$
@@ III S3E Dataset
The S3E dataset has been meticulously assembled, prioritizing high temporal precision in the synchronization of data from an array of sensors. Each sensor is calibrated to a shared timescale, ensuring that the captured multi-sensory information is harmonized. This precise synchronization is facilitated by an advanced time synchronization mechanism, essential for the accurate integration and co-registration of data across diverse modalities. For a visual representation of the sensor placement on our mobile robot platform, refer to LABEL:Fig:sensor_layout. Our mobile platforms are offered in two specialized versions, each tailored to meet specific operational needs:
  • S3Ev1.0 (LABEL:Fig:platform_v1.0): Customized for indoor operations, this version boasts a compact design that guarantees exceptional maneuverability, enabling it to traverse tight indoor spaces effortlessly.
  • S3Ev2.0 (LABEL:Fig:platform_v2.0): This enhanced version features an expanded width, accommodating a state-of-the-art, higher-precision ground truth system. The integration of an UWB module significantly bolsters the platform’s localization accuracy, particularly in environments where traditional SLAM systems may falter.
The subsequent sections delve into a comprehensive description of the S3E dataset’s components and their distinctive features, highlighting the meticulous approach taken to ensure the dataset’s utility and reliability in advancing SLAM research.
@@ III-A Sensor Configuration
(a) S3Ev1.0.
(b) S3Ev2.0.
Figure 3: Time Synchronization Scheme. The scheme addresses synchronization in both outdoor environments using GNSS time as the global reference and in GNSS-denied environments like indoors and tunnels, where agents rely on a PTPv2 server for time synchronization.
@@
For a comprehensive understanding of the data acquisition process and to facilitate replicability of our experiments, LABEL:Tab:hardware_properites outlines the technical specifications of the sensors and ground truth devices that were integrated into our payload. This includes the sensor types, their resolution, measurement range, accuracy, and any other pertinent technical details that define their contribution to the SLAM system’s performance.
@@ III-B Sensor Synchronization
Our synchronization system is built around an Altera EP4CE10 FPGA board acting as the primary trigger device, with an Intel NUC11TNKv7 serving as the host computer. The system is designed with a comprehensive set of I/O interfaces to accommodate the diverse requirements of different sensors, as depicted in LABEL:Fig:Time-Synchronization-Scheme..
For synchronization across agents, we address two distinct scenarios:
@@
Within each agent, the FPGA orchestrates the timing for various sensors:
  • S3Ev1.0: The FPGA periodically issues a pulse from a low-drift oscillator to trigger the LiDAR, stereo cameras, and IMU, as shown in LABEL:Fig:time_sync_v1.0. Specifically, the FPGA generates a 1 Hz pulse to initiate the LiDAR, which then outputs data at 10 Hz and updates its internal counter register upon receiving the trigger signal. The stereo cameras and IMU are designed to immediately return data following the receipt of an FPGA trigger pulse. The FPGA is configured to generate trigger pulses at 10 Hz for the cameras and 100 Hz for the IMU, tailored to the specific data output needs of these sensors.
  • S3Ev2.0: This version enhances synchronization by aligning the FPGA board with the pulse-per-second (PPS) signal from its internal GNSS/INS module, as depicted in LABEL:Fig:time_sync_v2.0. This alignment generates frequency trigger signals for IMU, stereo cameras, and LiDAR clock synchronization. The UWB receiver, not supporting trigger-based data acquisition, collects and outputs measured distance data directly.
@@
The calibration process was meticulously carried out in a series of five distinct stages, each designed to ensure the precise alignment and integration of sensor data:
Camera Intrinsic Calibration: We executed a high-precision intrinsic calibration for our cameras using the conventional checkerboard pattern, a method renowned for its reliability [27]. This process determined the camera intrinsic matrix $K$ and the lens distortion coefficients $D$, which are essential for correcting lens distortions and enhancing image accuracy.
IMU Intrinsic Calibration: Intrinsic calibration of the IMU involved accounting for white noise $\sigma_{n}$ and random walk noise $\sigma_{m}$ in the measurements. Building upon the nominal values provided in the Xsens MTi-30-2A8G4 IMU’s data sheet, we refined these parameters through a calibration process utilizing Allan variance, thereby achieving a more accurate probabilistic model of IMU noise [28].
Camera-Camera Extrinsic Calibration: The extrinsic calibration of the stereo cameras was performed to determine the rotation matrix $R$ and the translation vector $t$, which were computed by minimizing reprojection errors. This step is crucial for establishing the geometric relationship between the two cameras in a stereo setup [27].
Camera-IMU Extrinsic Calibration: The spatial transformation $T_{ic}$ between the camera and IMU was ascertained using the open-source Kalibr toolbox [29]. This calibration is essential for the integration of IMU data with visual data for enhanced pose estimation and sensor fusion.
Camera-LiDAR Extrinsic Calibration: The extrinsic calibration between the camera and LiDAR, represented as $T_{lc}$, was established by minimizing line-to-plane errors, ensuring the accurate co-registration of LiDAR point clouds with camera imagery [30].
@@ III-C Ground Truth
To establish accurate ground truth tracks for our dataset, we implement a multifaceted approach leveraging three distinct methods.
In outdoor settings where GNSS signals are readily accessible, we employ a dual-antenna Real-Time Kinematic (RTK) device to capture localization data with an accuracy reaching centimeter-level precision. For the S3Ev2.0 platform, we have made a significant advancement in our ground truth instrumentation by integrating a GNSS/INS unit. This upgrade enables high-frequency positioning data output, even during GNSS signal outages, such as those encountered in tunnel environments. The system achieves this by fusing RTK measurements with data from the built-in IMU through a post-processing fusion technique, thereby ensuring continuous and reliable ground truth data for our dataset.
In GNSS-denied environments, such as indoor laboratories, we employ a motion capture system equipped with 17 high-frequency cameras to record the start and endpoints of the tracks. This setup is utilized due to the prohibitive costs associated with constructing a large-scale motion capture system. For indoor scenarios not equipped with a motion capture system, such as in teaching buildings, we record the initial and terminal points of the tracks using the RTK device.

### v7

@@ I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is crucial for multi-agent cooperation, improving the robustness and efficiency of localization and mapping in shared spaces [1, 2, 3, 4]. Despite progress, repeatability and benchmarking in C-SLAM research face challenges due to: 1) System Complexity: C-SLAM systems’ intricate software and sensor integration make perfect replication difficult, highlighting the need for public datasets for methodological evaluation. 2) Evaluation Methodology: Traditional approaches, which divide single-agent SLAM datasets into segments for virtual agents, are impractical for real-world scenarios due to uniform perspectives and lighting at intersection points [5]. These limitations underscore the importance of developing more effective evaluation strategies and datasets to advance C-SLAM research.
@@
Furthermore, effective trajectory design for multi-robot scenarios in C-SLAM must follow two key principles: 1) Temporal and Spatial Diversity: Trajectories should be crafted to provide varying intra- and inter-loop closures at different times and places, ensuring robots have complementary perspectives. 2) Communication Constraints: Robots are typically limited to sharing information within close proximity, necessitating trajectory designs that maintain a reasonable interaction distance for effective communication. In this paper, we introduce four trajectory prototypes designed to meet these principles and evaluate the adaptability of C-SLAM methodologies across diverse closure strategies in multi-robot operations.
Publicly available datasets are crucial for the SLAM research community for two main reasons: 1) Development Acceleration: Conducting specialized SLAM experiments involves significant investment in advanced hardware and complex software, including calibration and ground truth generation. Datasets aid in streamlining this development process. 2) Benchmarking and Analysis: Datasets offer a standardized framework for comparing and assessing SLAM algorithms’ performance. While various SLAM datasets [6, 7] have been introduced over the years, contributing to sensor diversity and scenario breadth, most focus on single-agent systems. This focus may limit their utility for multi-agent scenarios, highlighting the need for more datasets that address multi-agent challenges.
The scarcity of datasets designed for C-SLAM is primarily due to the increasing challenges in data acquisition with more collaborating agents. Current C-SLAM datasets [8, 9, 10, 11, 12, 13] often focus on a single cooperative approach with significant spatial overlap, neglecting areas with minimal overlap. To fill this gap and enhance C-SLAM research, we introduce S3E dataset, offering a multimodal perspective with a variety of cooperative trajectory patterns in both outdoor and indoor environments.
Based on the S3E dataset, we conducted extensive experiments with cutting-edge SLAM methods across various applications, from single-robot to collaborative scenarios, using diverse sensor data. The findings reveal that current C-SLAM techniques struggle with inter-loop closures in some sequences, indicating a need for further research to enhance the robustness and reliability of these systems.
@@
  • We have created a cutting-edge C-SLAM dataset using three ground robots, each equipped with a 16-beam 3D laser scanner, two high-resolution color cameras, a 9-axis IMU, a UWB receiver, and a dual-antenna RTK receiver. This dataset is the first to incorporate UWB relative distance measurements, providing a new research dimension.
  • To assess C-SLAM’s performance in environments with limited overlap, we have captured extensive long-term sequences using four meticulously designed trajectory paradigms that address various intra- and inter-robot loop closure scenarios.
  • We have conducted a thorough and multifaceted evaluation of current C-SLAM techniques, employing both quantitative metrics and qualitative analyses to assess their performance and utility.
@@ II Related Work
Single-Agent SLAM Datasets. High-quality, extensive datasets are crucial for advancing SLAM research. The KITTI dataset [6], known for its comprehensive urban data, has been a significant benchmark for vision and LiDAR-based SLAM techniques. However, it has limitations, such as the absence of perfectly synchronized data, offering only estimated time delays between sensors. To address this, the EuRoC dataset [7] was introduced, offering synchronized visual and inertial data from a micro aerial vehicle in controlled settings, using hardware clocks for synchronization.
Multi-Agent SLAM Datasets. The UTIAS dataset [8] was a pioneering work with five robots exploring a small area. The AirMuseum dataset [10] introduced a heterogeneous multi-robot setup in a warehouse, and the FordAV dataset [9] presented a 3D map with multiple autonomous vehicles capturing seasonal changes. However, these datasets often have significant co-view overlap, limiting comprehensive C-SLAM evaluation.
The GRACO dataset [12] attempted to address this by introducing a ground-aerial approach, but it had synchronization issues and did not fully handle dynamic environments. The SubT-MRS dataset [13] added more variety with multi-robot SLAM data across different environments. Kimera-Multi [11] advanced the state with a comprehensive collection featuring diverse trajectory designs and multiple robots.
@@
It’s important to note that existing datasets, including GRACO, SubT-MRS, and Kimera-Multi, often overlook the complexities of collaborative trajectory patterns. The ability to handle intra- and inter-loop closures is crucial for C-SLAM robustness and is not consistently addressed. Considering these collaborative patterns is essential for developing C-SLAM algorithms that can operate effectively in environments with varying agent interactions and dynamic scenes.
@@
In this paper, we introduce a large-scale C-SLAM dataset that includes visual ambiguities and dynamic objects, captured under four distinct trajectory paradigms to assess C-SLAM generalizability under restricted overlap conditions. Our dataset also incorporates UWB relative measurement positioning, adding a significant dimension for C-SLAM research. For a comprehensive comparison of different datasets, including our contribution, refer to LABEL:Tab:dataset_comparison.
@@ III S3E Dataset
The S3E dataset is meticulously assembled with a focus on high temporal precision in sensor data synchronization. Each sensor is calibrated to a shared timescale for harmonized multi-sensory information, facilitated by an advanced time synchronization mechanism. This is crucial for accurate data integration across different modalities. The dataset features two mobile robot platform versions:
  • S3Ev1.0: Designed for indoor use with a compact design for exceptional maneuverability in tight spaces.
  • S3Ev2.0: Enhanced with a wider frame to accommodate a high-precision ground truth system and an UWB module for improved localization accuracy in challenging environments.
For a visual representation of sensor placement, refer to LABEL:Fig:sensor_layout.
@@ III-A Sensor Configuration
The technical specifications of the sensors and ground truth devices integrated into our payload are outlined on our project website [^1] https://pengyu-team.github.io/S3E. This includes the sensor types, their resolution, measurement range, accuracy, and any other pertinent technical details that define their contribution to the SLAM system’s performance.
@@ III-B1 Time Synchronization
Our synchronization system is built around an Altera EP4CE10 FPGA board acting as the primary trigger device, with an Intel NUC11TNKv7 serving as the host computer. The system is designed with a comprehensive set of I/O interfaces to accommodate the diverse requirements of different sensors. For synchronization across agents, we address two distinct scenarios:
@@ III-B2 Sensor Calibration
The calibration process was meticulously executed through five distinct stages: camera-intrinsic, IMU-intrinsic, camera-camera extrinsic, camera-IMU extrinsic, and camera-lidar extrinsic. Each stage was crucial for ensuring the accurate alignment and integration of sensor data.
@@ III-C Ground Truth
To create accurate ground truth tracks for our dataset, we use a three-pronged approach:
For outdoor environments with good GNSS signal reception, a dual-antenna RTK device is used to achieve highly accurate localization data with centimeter-level precision. The S3Ev2.0 platform features a GNSS/INS unit that significantly enhances ground truth instrumentation. This integration allows for high-frequency positioning data output even during GNSS signal outages, such as in tunnels. The system ensures continuous and reliable ground truth data by fusing RTK measurements with IMU data using post-processing fusion techniques.
In indoor environments without GNSS signals, like laboratories, a motion capture system with 17 high-frequency cameras is used to record track start and endpoints due to the high costs of a large-scale system. In other indoor areas lacking a motion capture system, such as teaching buildings, the RTK device is used to record the initial and terminal points of the tracks outside the buildings.